A Bibliography of Publications in ACM SIGGRAPH Computer Graphics: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.13

Title word cross-reference

+ [Wol00]. 3 [AC01, Ama08, BSLM01, Cai05, CBC^{+}01, Cav00, CB00, DG00, Hib02a, HSKK01, Jon03, KB00, Kli00, MMS00, Meg00, PvdDJ^{+}01, Par00b, Pin00, RV00, Rui00, Sch08a, Sha00b, TCMS04, Tho04a, Umb06, Wit07]. = [Wol00]. z [WFP^{+}01].

-buffer [WFP^{+}01].

16th [Ama08].

2 [SEP^{+}01]. 2000 [Fiu01a, Rhy00a, Rhy00b]. 2001 [Ano01a, Ano01c, Ano01b, Ano02d, Ell01]. 2002 [Ano02c, Ano02g, Ebe02, Ell02a]. 2003 [Bra04, Ebe03]. 2004 [GA04c]. 2008 [Bar08].

= [WLS00].

abstractions [Wes04d]. accepts [Slo06]. accessible [Meg00]. Achievement [Ano01a]. ACM [Ano01a, Ano01c, Ano01b, Gre06, Ano00a, Ano02c, Ano02d, Ano02e, CK05, Ell00c, GA04b, GA04a, Leo02b, Leo02a, Rhy00a, Tuf07, Woo04]. acoustics [TFNC01]. Active [Tat00]. activity [Ell01, NER01, WM03]. Adaptive [FFBG01, DDCB01]. Adaptively [Sul02]. adding [IYNK01]. addresses [Ell00a]. Adrenaline [Hoo01]. advanced [Lem01]. advantages [EW00b]. aesthetic [DTB02]. AFIS [Zir07]. Against [Cra02]. age [Dar04, Pet07, Wes03b]. agitating [Dia09]. ahead [Mac00a]. air [Zen04]. AirlinX [Jon03]. Alain [Fiu01a]. algorithm [Ost01, WFP^{+}01]. algorithms [Zir07]. Alice [PF00]. analogies [HJO^{+}01]. analysis [Hib03d]. Anatomical [RS09]. Anatomists [Cra02]. Ancient [Lem01]. Andrew [Ano01a]. Angeles [ACM01, Ano02d]. animal [KNB^{+}04]. animate [Kol08, PF00]. animated [BN06]. Animating [Awa01, Las01]. Animation [CVB01, Wol00, BE01, Cai05, Che01, CJ01, EB00, FvdPT01, FF01, Her04, MS01, MTAS01, Swi08, Tho04b, Wit07, WLS00, vdDKP01]. animator [Pit03]. animators [Gol01]. Announcements [Ano00b]. anything [Sim00a]. appearance [WM01]. Application [Ano02e, TDM00]. applied [Lem01]. Applying [WM01]. approach [SGH^{+}01, TPLD04]. Approximate [BKZ01]. arbitrary [WL01]. archeological [Lem01]. architectural [LDG01]. area [WM03]. Art [Pet07, Bes03, EW00a, EW00b, Ebe02, Ebe03, Huh04, Kal04, Wes03b]. Artificial [TPLD04]. Artist [Wes00c, Eas06, Wes00e, Wes02b]. Artists [Cra02, KM01]. arts [Ebe03, Wes09]. Asia [Wes07a]. Atlanta [Ano02d]. Attract [WEL02]. Audience [WEL02]. Augmented [Lar06, BGL05, Dia09]. August [ACM01, Ano02d]. automatic [CS01, HSKK01]. Automating [SM01]. Avatar [Lap00]. Avatars98 [DG00]. Avatars99 [DG00]. Award [Ano01a, Ano01c, Ano01b]. Awards [Ano01a, Ano01c, Ano01b].

baby [Ben06]. backto [Mac00a]. backward [CB00]. bandwidth [Wal00]. bandwidth-starved [Wal00]. Bangkok [Tuf07]. based [BG01, BE01, DTB02, FvdPT01, Hib00c, Jag00, JDD^{+}01, LDG01, MS01, OCE01, OCDD01, Var05, vdDKP01, vdHD08]. basis [CBC^{+}01]. be [Lem01]. BEAT [CVB01]. been [Sim00a]. beginning [CHLS04, Spe05a]. begins [Car00c]. behavior [PvdDJ^{+}01, CVB01]. behind [Bra04]. being [TCMS04]. Ben [Slo06]. better [Cam01]. between [Bes00]. bidirectional [LYS01]. bigger [War06]. binary [KB00]. biological [Hib03d]. birth [Rou01b]. blur [BE01]. Boeing [BK02]. BOF [Rhy00b]. bolts [Che01]. book [Ell00c]. Boolean [BKZ01]. borders [Lew09]. Boston [Gre06]. BRDF [KE09]. BRDFs [KC08, MAA01]. breakdown [Rou01a]. Breaking [Sch08a]. bridges [Wes09]. brushes [BSLM01]. buffer [WFP^{+}01]. Building [Hib02a, Wes09]. business [Par00b].

CA [ACM01, Ano02d]. Cailliau

[Sch08b, Sch08c]. Calendar [Ano02f]. Call [Ano02d, Ano02g]. camera [SGH^{+}01]. can [Lem01, Leo02a, Zir07]. Canadian [Woo04]. capture [RS09, Sco03b]. cardiac [MP00]. career [Wes04c]. Carnegie [Slo06]. Carto [Rhy00b]. cartoon [Fox00, Lar06]. castles [Ama08]. cellular [LDG01]. center [McA03, Rhy00c]. Century [Ama08]. CFP2000 [Ell00c]. chair [Slo06]. Challenges [Tal04]. chapter [CK05, Leo02b, Zha07, Tuf07]. Chapters [Ano00c, Ano02j, Ano02h, Ano02i, Ano04e, Ano05c, Lan00a, Cle00, Woo04]. character [DTB02, FvdPT01]. character-based [DTB02]. characters [Hib02b, Las01, PF01, Sim00b]. China [SP01]. cities [PM01]. cloning [yNN01]. closer [Sco03a]. clusters [HEB^{+}01, SEP^{+}01]. co [Sch08b, Sch08c]. co-developer [Sch08b, Sch08c]. cold [Bes00]. collaborating [KM01]. collaborative [EB00, Ma07a]. come [Kol08]. comes [Wes03b]. Comics [Lan00b, Lan01, Lan02a, Lan02b, Lan02c]. commentary [Eas06]. Committee [Ell00a, Ell02b, NER01, Ano00a, Ano00e, Ano02d]. Communication [Wes03a, McA03]. communicator [Sim00b]. community [Rhy00a, RV00]. complex [DDB00, WFP^{+}01]. Components [Hib02a]. Composable [FvdPT01]. compression [AD01, Zir07]. computational [Hib01b]. Computer [Ano01a, Ano01c, Ano02k, Ano02l, CESA01, CEK^{+}02, DH00, Mac00a, Mac00b, Mac01a, McA03, PaHL00, Rou00a, SP01, Wes03b, Wes03c, Wes05a, Wol00, BST^{+}00, Cam01, CJ01, CHLS04, EW00a, EW00b, EB00, Gol01, KE09, Las01, Mac00c, MP00, Rhy00a, Rou01a, Spe05b, Tho04b, TPLD04, Wes04a, Wes04b, Wes04c, Wes06a, Wes06c, WLS00, Wes06b]. computer-generated [Spe05b]. computing [YM04]. Concept [Gol01]. condition [WM03]. Conference [ACM01, Ano02d, Ano02m, Ano04f, CR00, Rhy00a, War06, Wes02c]. Confessions [Hib00a]. conflict [Hoo01]. Connecting [Lew09]. considerations [RS09]. Consistent [PSS01]. console [Rou01b]. Constrained [Lév01]. constructionist [TPLD04]. content [Lek02, Mil06b, Rou01a, Wal00]. continue [Lan00a]. continues [Ell02b]. continuing [Wes02a]. Contribution [Ano01c]. Contributions [Ano02n, Ano05d, Ano05e, Hib00b]. controllers [FvdPT01]. controversial [Ell00a]. conversion [CS01]. cool [Tho04a]. Coons [Ano01c]. could [Wil06]. course [CHLS04, EB00]. Cover [Sul00, Ano02a, Ano02b, Ano03a, Ano03b, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b]. Creating [Ma07a, Sco03a]. Creation [Urs04, Lek02]. Creative [Ano01c, Gol01, Wes02b]. creativity [Gol01]. crossing [Lew09]. crossroads [Huh04]. crucial [DKJ06]. curriculum [CJ01]. cyberconferences [DG00].

D [Cav00, CB00, AC01, Ama08, BSLM01, Cai05, CBC^{+}01, DG00, Hib02a, HSKK01, Jon03, KB00, Kli00, MMS00, Meg00, PvdDJ^{+}01, Par00b, Pin00, RV00, Sch08a, Sha00b, TCMS04, Tho04a, Umb06, Wit07]. DAB [BSLM01]. Daniel [Lin07]. data [Eic00, KBSS01, Ma07a, Slo06, Zir07]. Dawn [BB04]. Debevec [Ano01b]. decorative [Hau01]. Dedication [Fiu01a]. Deepwave [Car00a]. defining [Wes03c]. deformations [DDCB01]. demo [BB04]. depths [Wes06a]. design [Slo06]. Designing [BGL05]. desktop [Lek02]. detailed [NER01]. developer [Sch08b, Sch08c]. Developers [Rhy00a]. Developing [DTB02]. development [MMS00, Sau03]. devices [Wal00]. diary [Die00]. Diego [Bus07]. diffraction [TFNC01]. diffuse [BG01]. diffusion [Ost01]. dig [Lem01]. digital [Dar04, Ebe03, Ell00c, Jon03, PF01, Pet07, Wes00e]. dimensional [Eic00, MP00, Tup01]. direction [Ell02a]. discovery [Wes04a]. discussion [Ell00b, Ell02c]. display [KKYK01, SEP^{+}01]. Distance [Sul02]. dive [McC00]. do [Sim00a]. documenting [Gol01]. Doing [Gre05]. domain [Ell00c, Zir07]. dots [Lew09]. down [BGGH05]. Dr. [Sch08b, Sch08c]. dream [Mat00]. DWeb [Rui00]. Dynamic [DDCB01, Mil06a]. dynamics [MP00].

e-business [Par00b]. Early [BK02]. earth [HBSB02]. easier [Swi08]. easy [PF00]. easy-to-use [PF00]. editing [OCDD01]. Editor [GA02, Ano04g, Ano04h, Ano04i, Ano04j, Ano05f, Ano05g, Wes08a, Wes08b, Hib00b, Sul01]. Education [EW00a, Kol08, Wol00, CK05]. educators [BB01]. effective [AS01]. effects [Wil06, vdDKP01]. efficient [Ost01, RH01a]. efforts [Ell02b, Hib01c]. Eighth [Ano02o]. Einstein [Wes02a]. Elections [Ano00a, Ano02c]. electronic [KB00, McA03]. emerging [Ma04a]. emotional [Rou01a]. End [Lek02, Spe05a]. End-to-end [Lek02]. Energy [CK05]. engine [LKM01, Wil06]. engineering [Bro04]. Enormous [Wes00a]. entanglement [Mas04]. entry [ZCHS02]. environment [BST^{+}00, RH01a, RSR^{+}01]. environments [Rui00, Sco03a, TFNC01, Ano02o]. environments/ [Rui00]. EPA [Rhy00c]. error [Ost01]. error-diffusion [Ost01]. Escher [Jag00]. essentials [Hoo01]. estimation [HSKK01]. Eurographics [Ano02o]. events [DH00]. evolution [Dia09]. example [Zir07]. examples [Tay00]. excitement [Tal04]. Executive [Ano00e, Ano02d]. exhale [Sim00a]. exhibit [Bru00]. exhibition [Bar08]. experience [GA04a]. experiences [GA04b]. exploration [Wes04a]. Exploring [Ama08, Gil01]. Expression [CVB01, yNN01, LSZ01]. Expressive [LSZ01, LM02]. extends [Wes05b]. extraction [KBSS01]. extreme [HLM01]. eyes [Wes00a]. EZTV [Mas04].

facial [RS09]. factorization [MAA01]. Facts [Wes08b]. fairy [RV00]. fairy-tale [RV00]. fast [Sim00a]. Fear [Bus07]. Feature [KBSS01, LDG01]. Feature-based [LDG01]. FEELEX [IYNK01]. feet [Wes02c]. few [DH00]. fewest [Abt03]. fiction [Wes08b]. fidelity [WFM01b]. field [Hib01b, Ma04a]. Fields [Sul02]. files [Ano02t, Hib01b, Hib01c]. fills [Var05]. film [Umb06, Wit07]. filmmaking [Swi08]. Finding [Wes06b, Wes06c, Slo06]. first [War06]. fixed [Dia09]. fixed-mobile [Dia09]. flexible [BST^{+}00]. fluid [MP00]. Focus [Woo04]. focused [Eas06]. FoleyAutomatic [vdDKP01]. forecast [MMS00]. forest [Bru00, DDB00]. form [BKZ01]. formulae [Tup01]. formulation [BK02]. forwards [CB00]. found [Lem01].

Fournier [Fiu01a]. framework [RH01b]. free [BKZ01, Tup01]. free-form [BKZ01]. frustrations [KM01, Wes04b]. Fry [Slo06]. fuel [Tal04]. fully [HSKK01]. functions [CBC^{+}01, LYS01]. Funnelvision [Fox00]. future [BGGH05, Eds08, Umb06].

GA [Ano02d]. gait [SM01]. gallery [Ano00f, Ebe02, Ebe03]. Game [Wes05b, Wes04c, Rhy00a]. Games [Rou00a, Rou01a, Bla05, Dar07, Rhy00a, Swi08]. Gaming [Rou00a, Rou00b, Rou01b]. Gävle [ZCHS02]. Gearing [GA04c]. generalization [AS01]. generated [Spe05b]. generation [SM01]. geometry [PG01]. Germany [CESA01]. get [Lap00, Wes04d]. getting [Leo02a, Wit07]. Giloi [Mac01a]. global [MTAS01]. glossy [BG01]. good [GA04b, Hoo01]. Goodbye [Hib03a]. graduate [DKJ06]. Graphics [Ano01a, Ano01c, Ano02g, Bar00, Mac00b, Rou00a, Spa07, Wol00, Ano02k, Ano02l, Cam01, CESA01, CEK^{+}02, CHLS04, DH00, EW00a, EW00b, Eds08, HEB^{+}01, IYNK01, KE09, Mac00a, Mac00c, Mac01a, McA03, PaHL00, PF00, PMTH01, Rou00b, Rou01b, SP01, Tho04a, Tho04b, TPLD04, Wes03c, Wes04a, Wes04b, Wes04c, Wes05a, Wes06a, Wes06c, WLS00, Wes06b]. graphing [Tup01]. grasping [Wes00a]. Green [Mac00c]. grounds [Hau05]. grow [Lan00a]. growth [Wes03a]. Guerilla [Bra04]. Guest [Hib00b, Sul01]. guided [MTAS01]. gyromouse [DKJ06].

hands [Wes00a]. haptic

[BSLM01, IYNK01]. hardware [PMTH01]. hatching [PHWF01]. ’HDR [Mil06a]. heart [MP00]. Heat [WEL02]. heaven [Swi08]. hidden [Wes04d]. High [DKJ06, Mil06a, Cai05, MAA01, SGH^{+}01, SEP^{+}01, Zir07]. high-performance [MAA01, SEP^{+}01]. highly [WFP^{+}01]. historic [RF06]. history [Mac00c, Mac01b].

HIT [BGGH05]. Hokule’a [Wes02a]. Hollywood [Swi08]. Homomorphic [MAA01]. Hong [PaHL00]. Hopgood [Mac00c]. hostess [Par00a]. house [Mni00]. human [MP00, WM03, BGGH05]. humanities [McA03]. Hybrid [SGH^{+}01]. hydrological [TDM00]. Hypercosm [Meg00].

IBR [SGH^{+}01]. IceBorg [Bes00]. ideas [Spa07]. If [Kol08]. illumination [MTAS01]. Illustrative [RBGV08]. Image [BG01, BE01, EF01, HJO^{+}01, OCDD01, vdHD08, Bes03, SGH^{+}01]. Image-based [BG01, BE01, OCDD01]. Images [Ano02p, Wes00b, Wes00c, LSZ01, Wes03c]. imaginary [Ano00f]. imaging [KNB^{+}04]. immersive [RSR^{+}01, SR00]. improving [AS01]. individual [GA04b]. industry [Cal01, ZCHS02]. Information [WM03, Jon03, PMKL01, Rhy00c, Wes03a]. informed [Ell02c]. Innovations [Deb06]. insights [Roh07]. Installation [Bes03]. installations [DTB02]. instruction [Urs04]. Integrating [WFM01a]. Intel [Pit03]. intellectual [Ell00c]. intelligence [TPLD04]. intelligent [Hib01a]. interaction [PvdDJ^{+}01]. Interactive [KKYK01, Par00b, AC01, Bar00, BSLM01, DTB02, PF00, WFP^{+}01, vdDKP01]. Interactivity [LM02]. interdisciplinary [EB00]. Interface [BGGH05, Hib02a, DKJ06, Par00b]. interfaces [BGL05]. Internet [DG00]. Inventing [BGGH05]. inverse [RH01b]. investigation [BK02]. invisible [RF06]. irradiance [RH01a]. issue [Wes07b]. Issues [WEL02]. ities [Tho04b]. iVisit [Dia09].

Janimation [Her04]. Japanese [Ama08]. joys [Wes04b].

kelp [Bru00]. key [LM02]. Kizamu [PF01]. knitwear [XCL^{+}01]. knowledge [Ma07a].

Kong [PaHL00]. Korea [CEK^{+}02].

la fabrique [KB00]. Lab [BGGH05]. Laboratory [BGGH05, MMS00]. Lance [Ano01c]. landing [Kli00]. language [Bar00, Wes00d]. Large [Hib03b]. layout [Lem01]. Learning [Spe05b, Bro04, Sco03a, Var05]. less [Bla05]. Library [Hib02a]. lies [Cor09]. life [Lap00, Sco03b, Bar08]. light [JMLH01, WM01]. Lightning [SEP^{+}01]. Lightning-2 [SEP^{+}01]. like [Wes02a]. Liming [Zha07]. line [Ell00c, Mat00, RV00]. Lingafelter [Lin07]. liquids [FF01]. Listing [Ano00c, Ano02j]. literacy [EW00a, EW00b]. literal [Wes03c]. little [Teo07, Wil06]. loafers [Bus07]. long [Bra04, GA04a]. longer [Wes00b]. look [Bra04]. looking [CB00, Mac00a]. Lord [Sco03b]. lossless [AD01]. lot [Swi08]. lumigraph [BBM^{+}01]. lumislice [XCL^{+}01]. Lunar [Kli00].

machines [Hib01a]. made [Ell02a, Swi08]. MagicBlossom [Sha00a]. Making [Umb06, Wes02b, Wes06d, Meg00, Tho04a]. mammalian [WFM01a]. manifold [WL01]. MANs [KC08]. mapping [Gol01, Lév01, LSZ01, SSGH01]. maps [AS01, FFBG01, MGW01, RH01a]. match [Swi08]. matching [HSKK01]. mathematical [Tup01]. me [Ben06]. meaning [Wes03c]. measurement [CR00]. Measuring [WFM01b]. media [Wal00]. media-rich [Wal00]. medium [DG00, Kal04]. Meeting [Ano02d]. meets [Rhy00b]. Mellon [Slo06]. Members [Ano04k, Ano04l, Ano04m, Ano05h, Ano05i]. Membership [Ano02e]. mesh [PSS01]. meshes [AD01, Lév01, SSGH01]. Message [Owe06]. method [AC01, Sau03]. methods [Slo06, Tup01]. microbrowser [Sha00b]. mind [Gol01]. Minutes [Ano00e, Ano02d]. Miranda [Eas06]. mixed [GA04d]. mobile [Dia09]. model [DDB00, JDD^{+}01, JMLH01, MP00, PF00]. Modeling [Abt03, Dar07, TFNC01, Wol00, Cai05, OCDD01, PM01, PMKL01, TDM00, WLS00]. modelling [vdHD08]. models [KE09, LDG01, WFM01a, WM01]. moments [Hoo01]. mosaics [Hau01]. motion [BE01, OCE01, RS09]. mountain [Roh07]. movie [TCMS04]. Movies [Rou00a, Dar07]. much [Sco03a, War06, Wes02c]. multi [Eic00, RSR^{+}01]. multi-dimensional [Eic00]. multi-user [RSR^{+}01]. multimedia [Eas06]. museum [KB00, Mas04]. music [Car00a]. musical [RSR^{+}01]. Mystery [Mni00]. myth [Roh07].

names [Ell00c]. neat [Gre05]. NEC [Hib01c]. need [Var05]. nervous [Rou01a]. next [Cam00, Ma07b]. Nierenberg [Slo06]. night [JDD^{+}01]. no [Wes00b]. NOAA [MMS00]. Nobel [Wes01]. Nominating [Ano00a]. notes [Sco03b]. NURBS [BK02]. nuts [Che01]. NZ [BGGH05].

objects [CBC^{+}01, PvdDJ^{+}01]. October [Ano02d]. offering [GA04d]. oil [Hib03c]. old [Lem01]. on-line [Ell00c, Mat00, RV00]. one [Sch08b]. online [Mas04]. Open [Die00, Ell00b]. operated [McC00]. operational [Hib03b]. operations [BKZ01]. Oppor [Tho04b]. Oppor-toon-ities [Tho04b]. Opportunities [Mac00b, EW00b]. Optimization [MS01]. Optimization-based [MS01]. options [Gil01]. organization [Car00c]. ornamentation [Dar04]. Other [Lan00b, Lan02a, Ell00a, Lan01, Lan02b, Lan02c]. our [Cam01, GA04d]. outreach [Rhy00a]. Outstanding [Ano01c]. overview [SP01]. own [Hib03c].

Pacific [Ano02g]. paint [PF00]. Painting [Zen04, BSLM01]. Papers [Ano02g, Fiu01b]. paradox [Sch08a]. parameterizations [PSS01]. part [Sch08b, Sch08c, TCMS04]. participate [Cav00, Ell02c]. pattern [WFM01a]. Paul [Ano01b]. PC [Rou01b, SEP^{+}01]. Penn [Zha07]. people [DH00]. perceived [War06]. Perception [MTAS01, CR00, Lar06]. Perception-guided [MTAS01]. performance [Mat00, MAA01, Sco03b, SR00, SEP^{+}01]. personal [Pet07]. perspective [BN06, Pet07]. photo [OCDD01]. Photorealistic [XCL^{+}01]. physical [PvdDJ^{+}01]. physically [JDD^{+}01, OCE01, vdDKP01]. physically-based [JDD^{+}01, vdDKP01]. physics [FvdPT01]. physics-based [FvdPT01]. pioneering [DG00]. Pioneers [Mac00b, Ano02l, Mac00a, Mac01a]. pipeline [Car00c, Tho04a]. Pittsburgh [Abt03]. place [Bes00]. placemaking [Roh07]. places [DH00]. Plague [Wes00c, Wes00e]. plants [PMKL01]. play [DKJ06]. Playing [Wes04c]. Plenoptic [AC01]. point [PG01]. point-sampled [PG01]. Policy [ELS00, WEL02, Ano02q, Ano02r, Ano02s, Ell00b, Ell00c, Ell01, Ell02b, Ell02c, NER01]. polygonal [Lév01]. polygons [Abt03]. Polynomial [MGW01]. polynomiography [Kal04]. positional [PMKL01]. possible [Abt03, Wes00d]. Postprocessing [Wol00, WLS00]. Practical [FF01, Tay00, JMLH01]. pre [DKJ06]. pre-teens [DKJ06]. predicting [WFM01b]. Preface [Fiu01b]. prepares [ZCHS02]. preparing [Wes04c]. Preproduction [Cal01]. presence [Dia09]. presenter [Die00]. president [Owe06]. principles [Lem01]. Prizes [Wes01]. Procedural [PM01, PMTH01]. Proceedings [ACM01]. process [Ebe02, Gol01, Wes02b]. processing [PG01, RH01b]. produce [Zir07]. product [Ebe02]. production [Lek02]. Professional [Cle00]. programmable [LKM01, PMTH01]. programming [Meg00]. Progress [Ell02a]. Progressive [AD01, SSGH01]. Project [IYNK01]. property [Ell00c]. provides [Ell02a, Jon03]. Public [Ano02q, Ano02r, Ano02s, Ell00c, ELS00, Ell01, Ell02b, WEL02, Ell00b, Ell02c, NER01]. puzzle [Wes02a].

Quantum [Mas04]. quickly [Ma04a]. quilting [EF01].

radial [CBC^{+}01]. randomized [WFP^{+}01]. range [Mil06a]. ratio [LSZ01]. ratios [Zir07]. reach [Wes05b]. readers [GA04d]. Real [PHWF01, SR00, Ama08, DDCB01, Lar06, LYS01, PMTH01, Umb06, Wes06d]. Real-time [PHWF01, SR00, Ama08, DDCB01, Lar06, PMTH01]. real-world [LYS01]. reality [Bes03, BGL05, Dia09, DKJ06, Sau03]. reconstructing [AC01]. Reconstruction [CBC^{+}01]. refines [DKJ06]. reflection [WM01]. reflections [Lar06]. Reliable [Tup01]. Remembrances [Mac00b]. Remotely [McC00]. Rendering [AS01, Wol00, WLS00, BG01, BBM^{+}01, CR00, Deb06, Lar06, MAA01, MTAS01, RH01b, WFP^{+}01, XCL^{+}01]. Report [Ano00a, Ano02m, CR00, Lin07, Rhy00a, Zha07]. reports [Ano04f, CR00]. representation [CBC^{+}01, RH01a]. Representing [KC08]. research [Ma07b, McA03]. Researcher [Ano01b]. resolution [SGH^{+}01]. resources [EW00b]. responsibility [Mil06b]. Reversals [Wes00c, Ano02p, Wes00b]. revolution [Cam00]. Ricardo [Eas06]. rich [Wal00]. right [Wit07]. Rings [Sco03b]. Rivet [BST^{+}00]. roads [Hib03c]. Robert [Sch08b, Sch08c]. rock [Bes00]. role [DDB00, DKJ06]. Roomz [Str00]. roundUP [Cav00, CB00, Die00]. route [AS01]. ROV [McC00]. running [Sim00a].

Sampled [Sul02, PG01]. sampling [DDCB01]. samurai [Ama08]. San [Bus07]. scalable [AC01, HEB^{+}01]. Scanning [PvdDJ^{+}01]. scene [CS01]. scenes [Bra04, WFP^{+}01]. school [DKJ06, Mac01a]. schools [Cai05]. science [Bro04, CHLS04, EW00a, EW00b, HBSB02, Wes09]. Scientific [KNB^{+}04, Rhy00c, Tay00, YM04]. scrolls [Lem01]. sculpting [PF01]. Second [Bar08]. secret [Cal01]. security [Ma04b]. SEDRIS [Car00c]. see [Wes04d]. sensitive [KBSS01]. separated [Rou01b]. September [Ano02d]. sequences [SGH^{+}01]. Serious [Bla05]. sessions [Sco03b]. shading [PMTH01]. shadow [FFBG01]. shape [WFM01a]. shapes [HSKK01]. share [BB01, Ma07a, Mac00c]. shared [GA04b]. shipwrecks [RF06]. short [BN06]. Should [Hib03c]. show [Par00a]. Side [Lan00b, Lan02a, Lan01, Lan02b, Lan02c]. SIGGRAPH [Ano01a, Ano01c, Ano01b, Ano02c, Ano02d, Leo02a, Tuf07, Woo04, ACM01, Ano00a, Ano00e, Ano02e, Bar08, Bra04, CK05, Ebe02, Ebe03, Ell01, Ell02a, GA04b, GA04a, GA04c, Gre06, Leo02b, Lin07, NER01, Rhy00a, Rhy00b, Wes02c, Wes07a, Zha07]. SIGGRAPHICON [Eis07]. signal [RH01b]. signal-processing [RH01b]. Significant [Ano01b]. similarity [HSKK01]. simple [Ost01]. SimsalaGrimm [RV00]. Simulating [Hau01]. simulation [DDB00, FSJ01, TDM00, vdDKP01]. Singapore [Tuf07]. single [BG01]. skeptic [Hib00a]. Skin [BB04]. sky [JDD^{+}01]. small [KNB^{+}04]. smoke [FSJ01]. Social [Hib02b, Eas06]. solids [BKZ01]. solution [MTAS01]. solutions [Par00b]. SOMs [KC08]. sophomore [CJ01]. sound [vdDKP01]. sounds [OCE01]. space [DDCB01, Ma07a]. Spaces [Hib00d, Hib00e].

SpaceTime [Bar08]. Sparking [Sco03b]. Spatial [Car00b, WM03]. specific [Zir07]. Spectral [PG01]. specular [BG01]. splatting [ZPvBG01]. spurs [Ell01]. stage [Awa01, RSR^{+}01]. standard [Hib03c]. standardization [Car00b]. Standards [Car00c, WM01]. Standing [Huh04]. Starting [Leo02b]. starved [Wal00]. state [Ebe03]. Stay [Ell02c]. steering [SR00]. stepping [Rhy00c]. stereo [SGH^{+}01]. Stereoscopic [Wit07, KKYK01, SGH^{+}01, Sch08a, Umb06]. Steven [Ano01c]. Stirring [Sul00]. stitching [AC01]. stock [YM04]. stop [BE01]. Stopped [Wes00c, Wes00e]. storytelling [Bar00, Hoo01]. strange [Bra04]. Streaming [Pin00]. Student [Ano00f, BN06, CK05, Zha07]. students [EW00a, EW00b, ZCHS02]. Studio [Bra04]. studios [Lek02]. studying [MP00]. stuff [Gre05]. subsurface [JMLH01]. subsystem [SEP^{+}01]. success [BB01]. suck [Sim00a]. Sunny [Par00a]. Super [Hib01a]. Super-intelligent [Hib01a]. supercomputer [Hib01c]. Surface [ZPvBG01, IYNK01, KBSS01]. surfaces [BG01, LYS01, Tur01, WL01]. surge [Ma07b]. survey [Ell00c, KE09]. surveys [Hib00c]. SWAG [Ben06]. Sweden [Wil06]. symbolization [Spa07]. synthesis [EF01, SGH^{+}01, Tur01, WL01]. Synthesizing [LYS01, OCE01]. synthetic [Hib02b]. system [CS01, HBSB02, HEB^{+}01, Jon03, Lar06, PF01, PMTH01]. systems [BST^{+}00, MMS00].

Taking [YM04]. tale [RV00]. task [Ano00d]. tautology [RBGV08]. Teaching [Cai05, EW00b, Ma05, EW00a]. team [Hib00c]. TEATIME [Gre06]. techniques [Dar07]. Technology [BGGH05, Ano00d, Huh04, Lem01, Mil06a, RBGV08, Rhy00c, Umb06, Wes05b, Wes09]. teens [DKJ06]. tele [Dia09]. tele-presence [Dia09]. television [Eds08, Ell00c]. testing [HLM01]. text [CS01, Tho04a]. text-to-scene [CS01]. Texture [SSGH01, Tur01, WL01, EF01, Lév01, LYS01, MGW01]. texturing [LDG01]. themed [RV00]. theory [TFNC01]. There [Wes02c]. these [Lem01]. thing [Wes00b]. things [Teo07]. think [Sco03a]. thinkers [Wes01]. Thinking [Wes02a]. third [Ell00c]. Threading [CJ01]. three [KKYK01, MP00]. three-dimensional [MP00]. throughs [AC01]. time [Ama08, Cam00, DDCB01, Lar06, PHWF01, PMTH01, SR00]. tiny [Wes00a]. tired [Wes02c]. together [Urs04]. toiling [Rou00b]. tomorrow [Lem01]. too [Sim00a]. tool [Hib03d]. toolkit [HLM01, CVB01]. tools [Rou00b]. toon [Tho04b]. topics [Ell00a]. Topology [HSKK01]. Towers} [Sco03b]. toy [Jag00]. transfer [EF01]. transmission [AD01]. transport [JMLH01]. travelers [Jon03]. triangle [AD01]. Tricks [Las01]. Trip [Tuf07, Bra04]. twirl [Sha00b]. Two [Rhy00c, Sco03b, Hoo01, Sch08c, Tup01]. two-dimensional [Tup01].

U.S. [Rhy00c]. UK [DH00, Hib00c, Mac00c]. understand [Cam01]. understanding [DDB00]. uniform [TFNC01]. Univ [Zha07]. Universal [Wal00]. University [ZCHS02]. Unstructured [BBM^{+}01]. Update [ELS00, Cle00, CK05]. usability [AS01]. use [BK02, PF00, PMKL01]. useless [RBGV08]. User [Hib02a, Hib03b, LKM01, RSR^{+}01]. user-programmable [LKM01]. users [KKYK01]. Using [Cam01, Hib02a, DDCB01, KC08, TFNC01, XCL^{+}01].

value [Wes04d]. variables [Tup01]. vehicle [McC00]. verge [Rou01a]. versaCAD [Mac01b]. vertex [LKM01]. very [Bes00, SGH^{+}01, Zir07]. via [Lar06]. video [Lar06, Swi08, TCMS04]. video-to-cartoon [Lar06]. Videogames [Mil06b]. VideoTrace [vdHD08]. viewpoint [Rhy00c]. Virtual [Ano02o, Bru00, Par00a, Rui00, Sim00b, BSLM01, DKJ06, Hag00, Roh07, Sau03, Sco03a, TFNC01]. Vis [Ano02t, Hib01b, Hib01c]. VisAD [TDM00]. VisBio [Hib03d]. Visfiles [Ano02u, Ano04o, Ano05j, Ano05k, Hib00c, Hib00d, Ano02v, Ano03c, Ano04n, Ano04p, Ano04q, Ano03d, Ano03e, Ma07b, RBGV08]. visible [Wes02b]. Visual [Bro04, FSJ01, Hib00d, Wes00c, Wes01, Wil06, EW00a, EW00b, Kal04, WFM01b, Wes02a]. Visualising [RF06]. Visualization [Hib00e, HLM01, Hib02a, HBSB02, Ma04a, Ma04b, BST^{+}00, Car00a, DDB00, Hau05, Hib00a, Hib00c, Hib01b, Hib01c, Hib03b, Hib03d, KNB^{+}04, LM02, Ma05, Ma07a, Ma07b, McC00, MMS00, RBGV08, Rhy00c, SR00, Spe05b, Tay00, WM03, Wes00b, Wes00d, YM04]. Visualizing [Cor09, Eic00, Slo06]. volume [KBSS01]. VR2Go^{T M} [Sau03]. VRML [Mat00]. vs [Dar07].

walk [AC01]. was [Wes00c, Wes00e]. way [Wes06b, Wes06c]. Web [Cam00, Cav00, CB00, Fox00, Hib00c, Jag00, KB00, Kli00, PF00, Pin00, Var05]. Web-based [Hib00c, Jag00, Var05]. Web3D [WM03]. welcoming [Jon03]. we’ve [Sim00a]. whereas [Sim00a]. will [Kol08]. Williams [Ano01c]. WireGL [HEB^{+}01]. Witkin [Ano01a]. WordsEye [CS01]. work [Ano04k, Ano04l, Ano04m, Ano05h, Ano05i, Car00c, Tal04, Urs04]. Workshop [Ano02o, CR00]. World [Wes00c, Ama08, Ano02k, Cam01, LYS01, PaHL00, SP01, Wes00e]. worlds [Ano00f, Die00, Hag00, Tat00]. worth [Hag00]. WWW [Sch08b, Sch08c].

year [GA04a]. year-long [GA04a]. years [KB00].

Zealand [BGGH05]. Zhao [Zha07]. Zúñiga [Eas06].

References

Abt:2003:MPF

[Abt03] Kevin Abt. Modeling Pittsburgh … with the fewest possible polygons. Computer Graphics, 37(1): 25–26, February 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Aliaga:2001:PSS

[AC01] Daniel G. Aliaga and Ingrid Carlbom. Plenoptic stitching: a scalable method for reconstructing 3D interactive walk throughs. In ACM [ACM01], pages 443– 450. ISBN 1-58113-292-1. LCCN ????

ACM:2001:SCP

[ACM01] ACM, editor. SIGGRAPH 2001 Conference Proceedings, August 12–17, 2001, Los Angeles, CA. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-292-1. LCCN ????

Alliez:2001:PCL

[AD01] Pierre Alliez and Mathieu Desbrun. Progressive compression for lossless transmission of triangle meshes. In ACM [ACM01], pages 195–202. ISBN 1-58113-292-1. LCCN ????

Amakawa:2008:EWC

[Ama08] Jon Amakawa. Exploring the world of 16th Century Japanese castles and samurai in real-time 3D. Computer Graphics, 42(3):

3:1–3:??, August 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2000:ASE

[Ano00a] Anonymous. ACM SIGGRAPH 2001 elections: Nominating Committee report. Computer Graphics, 34(4):25, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2000:A

[Ano00b] Anonymous. Announcements.

Computer Graphics, 34(4):29– 30, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2000:CL

[Ano00c] Anonymous. Chapters listing.

Computer Graphics, 34(4):13– 14, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2000:TT

[Ano00d] Anonymous. Is the technology up to the task? Computer Graphics, 34(3):3–4, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2000:SEC

[Ano00e] Anonymous. SIGGRAPH Executive Committee minutes. Computer Graphics, 34(4):26– 27, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2000:SGI

[Ano00f] Anonymous. Student gallery:

imaginary worlds. Computer Graphics, 34(1):27–29, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2001:ASAa

[Ano01a] Anonymous. 2001 ACM SIG-GRAPH Awards — Computer Graphics Achievement Award — Andrew Witkin. In ACM [ACM01], page 11. ISBN 1-58113-292-1. LCCN ????

Anonymous:2001:ASAc

[Ano01b] Anonymous. 2001 ACM SIG-GRAPH Awards — Significant New Researcher Award — Paul E. Debevec. In ACM [ACM01], page 13. ISBN 1-58113-292-1. LCCN ????

Anonymous:2001:ASAb

[Ano01c] Anonymous. 2001 ACM SIG-GRAPH Awards — Steven A. Coons Award for Outstanding Creative Contribution to Computer Graphics — Lance Williams. In ACM [ACM01], page 12. ISBN 1-58113-292-1. LCCN ????

Anonymous:2002:ACa

[Ano02a] Anonymous. About the cover. Computer Graphics, 36(3):??, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:ACb

[Ano02b] Anonymous. About the cover. Computer Graphics, 36(4):??, November 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:ASEa

[Ano02c] Anonymous. ACM SIGGRAPH 2002 elections. Computer Graphics, 36(1):17–31, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:ASEb

[Ano02d] Anonymous. ACM SIGGRAPH Executive Committee minutes: Meeting October 28, 2001, Atlanta, GA; conference call, September 19, 2001; meeting August 18, 2001, Los Angeles, CA. Computer Graphics, 36(1): 32–35, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:ASM

[Ano02e] Anonymous. ACM SIGGRAPH membership application. Computer Graphics, 36(1):39, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:C

[Ano02f] Anonymous. Calendar. Computer Graphics, 36(1):37, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:CPP

[Ano02g] Anonymous. Call for papers:

Pacific Graphics 2002. Computer Graphics, 36(1):38, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:Ca

[Ano02h] Anonymous. Chapters. Computer Graphics, 36(2):??, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:Cc

[Ano02i] Anonymous. Chapters. Computer Graphics, 36(4):??, November 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:CL

[Ano02j] Anonymous. Chapters listing. Computer Graphics, 36(1):11– 12, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:CGA

[Ano02k] Anonymous. Computer graphics around the world. Computer Graphics, 36(3):??, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:CGP

[Ano02l] Anonymous. Computer graphics pioneers. Computer Graphics, 36(3):??, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:CR

[Ano02m] Anonymous. Conference report. Computer Graphics, 36 (3):??, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:Cb

[Ano02n] Anonymous. Contributions. Computer Graphics, 36(2):??, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:EEW

[Ano02o] Anonymous. Eighth Eurographics Workshop on Virtual Environments. Computer Graphics, 36(1):37, February 2002. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:IR

[Ano02p] Anonymous. Images and reversals. Computer Graphics, 36 (3):??, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:PPa

[Ano02q] Anonymous. Public policy.

Computer Graphics, 36(2):??, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:PPb

[Ano02r] Anonymous. Public policy. Computer Graphics, 36(3):??, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:PPc

[Ano02s]  Anonymous.  Public policy.

Computer Graphics, 36(4):??, November 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:VF

[Ano02t] Anonymous. Vis files. Computer Graphics, 36(2):??, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:Va

[Ano02u] Anonymous. Visfiles. Computer Graphics, 36(3):??, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2002:Vb

[Ano02v] Anonymous. VISFILES. Computer Graphics, 36(4):??, November 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2003:ACa

[Ano03a] Anonymous. About the cover. Computer Graphics, 37(2):??, May 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2003:ACb

[Ano03b] Anonymous. About the cover. Computer Graphics, 37(3):??, August 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2003:Va

[Ano03c] Anonymous. VISFILES. Computer Graphics, 37(1):??, February 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2003:Vb

[Ano03d] Anonymous. VisFiles. Computer Graphics, 37(2):??, May 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2003:Vc

[Ano03e] Anonymous. VisFiles. Computer Graphics, 37(3):??, August 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:ACa

[Ano04a] Anonymous. About the cover. Computer Graphics, 38(1):??, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:ACb

[Ano04b] Anonymous. About the cover. Computer Graphics, 38(2):??, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:ACc

[Ano04c] Anonymous. About the cover.

Computer Graphics, 38(3):??, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:ACd

[Ano04d] Anonymous. About the cover.

Computer Graphics, 38(4):??, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:C

[Ano04e] Anonymous. Chapters. Computer Graphics, 38(4):??, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:CR

[Ano04f] Anonymous. Conference reports. Computer Graphics, 38 (4):??, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Ea

[Ano04g] Anonymous. From the editor. Computer Graphics, 38(1): ??, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Eb

[Ano04h] Anonymous. From the editor. Computer Graphics, 38 (2):??, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Ec

[Ano04i] Anonymous. From the editor. Computer Graphics, 38 (3):??, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Ed

[Ano04j] Anonymous. From the editor. Computer Graphics, 38(4): ??, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:MWa

[Ano04k] Anonymous. Members at work. Computer Graphics, 38(2):??, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:MWb

[Ano04l] Anonymous. Members at work. Computer Graphics, 38(3):??, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:MWc

[Ano04m] Anonymous. Members at work. Computer Graphics, 38(4):??, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Va

[Ano04n] Anonymous. VISFILES. Computer Graphics, 38(1):??, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Vb

[Ano04o] Anonymous. Visfiles. Computer Graphics, 38(2):??, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Vc

[Ano04p] Anonymous. VISFILES. Computer Graphics, 38(3):??, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2004:Vd

[Ano04q] Anonymous. VISFILES. Computer Graphics, 38(4):??, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:ACa

[Ano05a] Anonymous. About the cover.

Computer Graphics, 39(1):??, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:ACb

[Ano05b] Anonymous. About the cover.

Computer Graphics, 39(2):??, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:Ca

[Ano05c] Anonymous. Chapters. Computer Graphics, 39(2):??, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:C

[Ano05d] Anonymous. Contributions.

Computer Graphics, 39(1):??, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:Cb

[Ano05e] Anonymous. Contributions.

Computer Graphics, 39(2):??, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:Ea

[Ano05f] Anonymous. From the editor. Computer Graphics, 39(1):

??, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:Ec

[Ano05g] Anonymous. From the editor. Computer Graphics, 39 (2):??, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:MWa

[Ano05h] Anonymous. Members at work. Computer Graphics, 39 (1):??, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:MWb

[Ano05i] Anonymous. Members at work. Computer Graphics, 39(2):??, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:Va

[Ano05j] Anonymous. Visfiles. Computer Graphics, 39(1):??, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Anonymous:2005:Vb

[Ano05k] Anonymous. Visfiles. Computer Graphics, 39(2):??, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Agrawala:2001:RER

[AS01] Maneesh Agrawala and Chris Stolte. Rendering effective route maps: improving usability through generalization. In ACM [ACM01], pages 241–249. ISBN 1-58113-292-1. LCCN ????

Awada:2001:AS

[Awa01]  Jef Awada.  Animating on

stage. Computer Graphics, 35 (2):42–44, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Barrett:2000:GLI

[Bar00] Mark Barrett. Graphics: the language of interactive storytelling. Computer Graphics, 34(3):7–10, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Barry:2008:SSS

[Bar08] Rick Barry. SIGGRAPH 2008 SpaceTime — Second Life exhibition. Computer Graphics, 42 (3):8:1–8:??, August 2008. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Belcher:2001:ESS

[BB01] John W. Belcher and R. Mark Bessette. MIT educators share success. Computer Graphics, 35 (1):18–21, February 2001. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Beeson:2004:SDD

[BB04] Curtis Beeson and Kevin Bjorke.

Skin in the “Dawn” demo. Computer Graphics, 38(2):14–19, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Buehler:2001:ULR

[BBM^{+}01] Chris Buehler, Michael Bosse, Leonard McMillan, Steven Gortler, and Michael Cohen. Unstructured lumigraph rendering. In ACM [ACM01], pages 425–432. ISBN 1-58113-292-1. LCCN ????

Brostow:2001:IBM

[BE01] Gabriel J. Brostow and Irfan Essa. Image-based motion blur for stop motion animation. In ACM [ACM01], pages 561–566. ISBN 1-58113-292-1. LCCN ????

Bendis:2006:SMB

[Ben06] Jared Bendis. SWAG me, baby! Computer Graphics, 40(3):10:1– 10:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Best:2000:IBR

[Bes00]  Andy Best.  IceBorg: be-

tween a rock and a very cold place. Computer Graphics, 34 (2):44–45, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Bestor:2003:IAI

[Bes03] Charles Bestor. Installation art: image and reality. Computer Graphics, 37(1):16–18, February 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Boivin:2001:IBR

[BG01] Samuel Boivin and André Gagalowicz. Image-based rendering of diffuse, specular and glossy surfaces from a single image. In ACM [ACM01], pages 107–116. ISBN 1-58113-292-1. LCCN ????

Billinghurst:2005:IFU

[BGGH05] Mark Billinghurst, Raphaël Grasset, Richard Green, and Michael Haller. Inventing the future down under: the Human Interface Technology Laboratory New Zealand [HIT Lab NZ]. Computer Graphics, 39 (2):18–23, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Billinghurst:2005:DAR

[BGL05] Mark Billinghurst, Raphael Grasset, and Julian Looser. Designing augmented reality interfaces. Computer Graphics, 39(1): 17–22, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Blomgren:2002:EIF

[BK02] Robert M. Blomgren and David J.

Kasik. Early investigation, formulation and use of NURBS at Boeing. Computer Graphics, 36 (3):27–32, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Biermann:2001:ABO

[BKZ01] Henning Biermann, Daniel Kristjansson, and Denis Zorin. Approximate Boolean operations on free-form solids. In ACM [ACM01], pages 185–194. ISBN 1-58113-292-1. LCCN ????

Blackman:2005:SGL

[Bla05] Sue Blackman. Serious games … and less! Computer Graphics, 39 (1):12–16, February 2005. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Bannan:2006:ASS

[BN06] Patrick Bannan and Thomas Netzband. The animated short: a student’s perspective. Computer Graphics, 40(2):4:1–4:??, May 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Braccio:2004:BSL

[Bra04] Peter Braccio. A behind the scenes look at The Guerilla Studio: the long, strange trip to SIGGRAPH 2003. Computer Graphics, 38(1):12–21, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Brown:2004:VLS

[Bro04] Judy Brown. Visual learning in science and engineering. Computer Graphics, 38(4):22– 23, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Brutzman:2000:VKF

[Bru00] Don Brutzman. Virtual kelp forest exhibit. Computer Graphics, 34(2):48–49, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Baxter:2001:DIH

[BSLM01] Bill Baxter, Vincent Scheib, Ming C. Lin, and Dinesh Manocha. DAB: interactive haptic painting with 3D virtual brushes. In ACM [ACM01], pages 461–468. ISBN 1-58113-292-1. LCCN ????

Bosch:2000:RFE

[BST^{+}00] Robert Bosch, Chris Stolte, Diane Tang, John Gerth, Mendel Rosenblum, and Pat Hanrahan. Rivet: a flexible environment for computer systems visualization. Computer Graphics, 34(1): 68–73, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Busse:2007:FLS

[Bus07] Mark Busse. Fear and loafers in San Diego. Computer Graphics, 41(4):1–3, November 2007. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cain:2005:TMA

[Cai05] Robin Cain. Teaching 3D modeling and animation in the high schools. Computer Graphics, 39 (2):24–25, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Caldwell:2001:PIS

[Cal01]  Craig Caldwell.  Preproduc-

tion: the industry secret. Computer Graphics, 35(2):30–31, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cameron:2000:TNW

[Cam00] Gordon Cameron. It’s time for the next Web revolution. Computer Graphics, 34(2):4–5, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cameron:2001:UCG

[Cam01] Gordon Cameron. Using computer graphics to better understand our world. Computer Graphics, 35(1):3, February 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Carroll:2000:DMV

[Car00a] Shane Carroll. Deepwave: music visualization. Computer Graphics, 34(2):47, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-

Carson:2000:SS

[Car00b] George S. Carson. Spatial standardization. Computer Graphics, 34(3):38–41, August 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Carson:2000:SPW

[Car00c] George S. Carson. Standards pipeline: work with the SEDRIS organization begins. Computer Graphics, 34(1):35–36, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cavallar:2000:WPW

[Cav00] Claudia Cavallar. Why participate in the Web3D roundUP? Computer Graphics, 34(2):63– 64, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Childs:2000:WRL

[CB00] Timothy Childs and Don Brutzman. Web3D roundUP: looking backward and forwards. Computer Graphics, 34(2):35–36, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Carr:2001:RRO

[CBC^{+}01] J. C. Carr, R. K. Beatson, J. B. Cherrie, T. J. Mitchell, W. R. Fright, B. C. McCallum, and T. R. Evans. Reconstruction and representation of 3D objects with radial basis functions. In ACM [ACM01], pages 67–76. ISBN 1-58113-292-1. LCCN ????

Chesnais:2002:CGK

[CEK^{+}02] Alain Chesnais, Jose Encarnação, Chang-Hun Kim, Myoung-Hee Kim, and Myung-Soo Kim. Computer graphics in Korea. Computer Graphics, 36(3):22– 26, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Chesnais:2001:CGG

[CESA01] Alain Chesnais, Jose Encarnação, Dietmar Saupe, and Marc Alexa. Computer graphics in Germany. Computer Graphics, 35(3):14– 21, August 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cheetham:2001:NBA

[Che01] Ed J. Cheetham. The nuts and bolts of animation. Computer Graphics, 35(2):48–52, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cunningham:2004:BCG

[CHLS04] Steve Cunningham, Werner Hansmann, Cary Laxer, and Jiaoying Shi. The beginning computer graphics course in computer science. Computer Graphics, 38(4):24–25, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cumbie-Jones:2001:TSC

[CJ01] Claudia Cumbie-Jones. Threading a sophomore computer animation curriculum. Computer Graphics, 35(2):53–56, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cohen:2005:EEA

[CK05] Adam Cohen and Matthew Kapfhammer. Energy and education: an ACM SIGGRAPH student chapter update. Computer Graphics, 39(2):13, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Cleary:2000:PCP

[Cle00] Colleen Cleary. Professional chapters: professional chapters update. Computer Graphics, 34 (1):20–21, February 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Correa:2009:VWL

[Cor09] Carlos D. Correa. Visualizing what lies inside. Computer Graphics, 43(2):5:1–5:??, May 2009. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Corson-Rikert:2000:CRR

[CR00] Jonathan Corson-Rikert. Conference reports: report of the workshop on rendering, perception, and measurement. Computer Graphics, 34(1):37–39, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Crawford:2002:AAA

[Cra02] Chris Crawford. Artists against anatomists. Computer Graphics, 36(1):8–10, February 2002. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Coyne:2001:WAT

[CS01] Bob Coyne and Richard Sproat. WordsEye: an automatic textto-scene conversion system. In ACM [ACM01], pages 487–496. ISBN 1-58113-292-1. LCCN ????

Cassell:2001:BBE

[CVB01] Justine Cassell, Hannes Högni Vilhjálmsson, and Timothy Bickmore. BEAT: the Behavior Expression Animation Toolkit. In ACM [ACM01], pages 477– 486. ISBN 1-58113-292-1. LCCN ????

Dar:2004:ADO

[Dar04] Vishal Dar. In the age of digital ornamentation. Computer Graphics, 38(3):24, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Dargie:2007:MTM

[Dar07] James Dargie. Modeling techniques: movies vs. games. Computer Graphics, 41(2):2:1–2:??, May 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Deutschman:2000:RVU

[DDB00] Douglas H. Deutschman, Catherine Devine, and Linda A. Buttel. The role of visualization in understanding a complex forest simulation model. Computer Graphics, 34(1):51–55, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Debunne:2001:DRT

[DDCB01] Gilles Debunne, Mathieu Desbrun, Marie-Paule Cani, and Alan H. Barr. Dynamic realtime deformations using space & time adaptive sampling. In ACM [ACM01], pages 31–36. ISBN 1-58113-292-1. LCCN ????

Debattista:2006:IR

[Deb06] Kurt Debattista. Innovations in rendering. Computer Graphics, 40(3):9:1–9:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Damer:2000:AAP

[DG00] Bruce Damer and Stuart Gold. Avatars98 and Avatars99: pioneering the medium of 3D Internet cyberconferences. Computer Graphics, 34(2):42, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Duce:2000:CGU

[DH00] D. A. Duce and F. R. A. Hopgood. Computer graphics in the UK: a few people, places and events. Computer Graphics, 34 (3):30–33, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Dialameh:2009:IAE

[Dia09] Orang Dialameh. iVisit agitating for the evolution of fixed-mobile tele-presence and augmented reality. Computer Graphics, 43(2): 3:1–3:??, May 2009. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Diefenbach:2000:OWD

[Die00] Paul Diefenbach. Open worlds: diary of roundUP presenter. Computer Graphics, 34(2):46, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Duncan:2006:HSG

[DKJ06] Matt Duncan, Matthew Kelley, and Jeffrey Jacobson. High school graduate refines gyromouse interface for virtual reality: pre-teens play crucial role. Computer Graphics, 40(2): 2:1–2:??, May 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Downie:2002:DAC

[DTB02] Marc Downie, Bill Tomlinson, and Bruce Blumberg. Developing an aesthetic: character-based interactive installations. Computer Graphics, 36(2):33–36, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Easterly:2006:RMN

[Eas06] Douglas Easterly. Ricardo Miranda Zúñiga: a multimedia artist focused on social commentary. Computer Graphics, 40(1): 4:1–4:??, February 2006. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ebert:2000:CIC

[EB00] David S. Ebert and Dan Bailey. A collaborative and interdisciplinary computer animation course. Computer Graphics, 34 (3):22–26, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Eber:2002:SAG

[Ebe02]  Dena Elisabeth Eber.  The

SIGGRAPH 2002 art gallery: process and product. Computer Graphics, 36(3):33–39, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Eber:2003:SAG

[Ebe03] Dena Elisabeth Eber. The SIG-GRAPH 2003 art gallery: state of art in digital arts. Computer Graphics, 37(3):20–28, August 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Edsall:2008:FTG

[Eds08]  Samuel Edsall.  The future

of television graphics. Computer Graphics, 42(2):4:1–4:??, May 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Efros:2001:IQT

[EF01] Alexei A. Efros and William T. Freeman. Image quilting for texture synthesis and transfer. In ACM [ACM01], pages 341–346. ISBN 1-58113-292-1. LCCN ????

Eick:2000:VMD

[Eic00]  Stephen G. Eick.  Visual-

izing multi-dimensional data. Computer Graphics, 34(1):61– 67, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Eisenberg:2007:S

[Eis07] Sharon Eisenberg. SIGGRAPH-ICON. Computer Graphics, 41 (4):1–2, November 2007. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2000:CAC

[Ell00a] Bob Ellis. Committee addresses controversial subjects… and other topics. Computer Graphics, 34 (2):31–33, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2000:ODP

[Ell00b] Bob Ellis. An open discussion on public policy. Computer Graphics, 34(3):34–37, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2000:PPD

[Ell00c] Bob Ellis. Public policy: domain names, digital television, CFP2000, ACM book on intellectual property and third on-line survey. Computer Graphics, 34 (1):33–34, February 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2001:PPS

[Ell01] Bob Ellis. Public policy: SIG-GRAPH 2001 spurs activity. Computer Graphics, 35(4):17– 20, November 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2002:PMS

[Ell02a] Bob Ellis. Progress made at SIG-GRAPH 2002 provides direction for 2003. Computer Graphics, 36 (4):15–19, November 2002. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2002:PPC

[Ell02b] Bob Ellis. Public policy committee continues efforts. Computer Graphics, 36(3):16–21, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2002:SIP

[Ell02c] Bob Ellis. Stay informed: participate in public policy discussion. Computer Graphics, 36 (2):13–22, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ellis:2000:PPP

[ELS00] Bob Ellis, Myles Losch, and Eugen H. Spafford. Public policy: Public policy update. Computer Graphics, 34(4):21–23, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Eber:2000:ETC

[EW00a] Dena Eber and Rosalee Wolfe. Education: teaching computer graphics visual literacy to art and computer science students. Computer Graphics, 34(1):24– 26, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Eber:2000:TCG

[EW00b] Dena Eber and Rosalee Wolfe. Teaching computer graphics visual literacy to art and computer science students: advantages, resources and opportunities. Computer Graphics, 34 (2):22–24, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Foster:2001:PAL

[FF01] Nick Foster and Ronald Fedkiw. Practical animation of liquids. In ACM [ACM01], pages 23–30. ISBN 1-58113-292-1. LCCN ????

Fernando:2001:ASM

[FFBG01] Randima Fernando, Sebastian Fernandez, Kavita Bala, and Donald P. Greenberg. Adaptive shadow maps. In ACM [ACM01], pages 387–390. ISBN 1-58113-292-1. LCCN ????

Fiume:2001:DAF

[Fiu01a] Eugene Fiume. Dedication — Alain Fournier, 1943–2000. In ACM [ACM01], page 10. ISBN 1-58113-292-1. LCCN ????

Fiume:2001:PP

[Fiu01b] Eugene Fiume. Papers preface.

In ACM [ACM01], page 9. ISBN 1-58113-292-1. LCCN ????

Fox:2000:FWC

[Fox00] Barrett Fox. Funnelvision Web cartoon. Computer Graphics, 34 (2):60–61, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Fedkiw:2001:VSS

[FSJ01] Ronald Fedkiw, Jos Stam, and Henrik Wann Jensen. Visual simulation of smoke. In ACM [ACM01], pages 15–22. ISBN 1-58113-292-1. LCCN ????

Faloutsos:2001:CCP

[FvdPT01] Petros Faloutsos, Michiel van de Panne, and Demetri Terzopoulos. Composable controllers for physics-based character animation. In ACM [ACM01], pages 251–260. ISBN 1-58113-292-1. LCCN ????

Galvis-Assmus:2002:E

[GA02] Patricia Galvis-Assmus. From the editor. Computer Graphics, 36(1):2, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-

Galvis-Assmus:2004:ASY

[GA04a] Patricia Galvis-Assmus. ACM SIGGRAPH: a year-long experience. Computer Graphics, 38 (3):2, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Galvis-Assmus:2004:ASI

[GA04b] Patricia Galvis-Assmus. ACM SIGGRAPH: individual experiences shared for the good of all. Computer Graphics, 38(4): 2, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Galvis-Assmus:2004:GS

[GA04c] Patricia Galvis-Assmus. Gearing up for SIGGRAPH 2004. Computer Graphics, 38(2):2, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930. Galvis-Assmus:2004:MOO

[GA04d] Patricia Galvis-Assmus. A mixed offering for our readers. Computer Graphics, 38(1):2, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Gilbert:2001:EO

[Gil01] Wayne Gilbert. Exploring options. Computer Graphics, 35 (2):38–39, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Goldstein:2001:CMM

[Gol01] Julie Goldstein. Concept mapping, mind mapping and creativity: documenting the creative process for computer animators. Computer Graphics, 35 (2):32–35, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Greenfield:2005:DNS

[Gre05] Gary R. Greenfield. Doing “neat stuff”. Computer Graphics, 39 (2):9–12, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Grey:2006:TBA

[Gre06] Jen Grey. TEATIME at Boston ACM SIGGRAPH 2006. Computer Graphics, 40(3):5:1–5:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hague:2000:VWW

[Hag00] James Hague. Are virtual worlds worth it? Computer Graphics, 34(2):7–10, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hausner:2001:SDM

[Hau01] Alejo Hausner. Simulating decorative mosaics. In ACM [ACM01], pages 573–580. ISBN 1-58113-292-1. LCCN ????

Hauser:2005:TNG

[Hau05] Helwig Hauser. Toward new grounds in visualization. Computer Graphics, 39(2):5–8, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2002:VES

[HBSB02] Bill Hibbard, Michael Böttinger, Martin Schultz, and Joachim Biercamp. Visualization in earth system science. Computer Graphics, 36(4):5–9, November 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Humphreys:2001:WSG

[HEB^{+}01] Greg Humphreys, Matthew Eldridge, Ian Buck, Gordan Stoll, Matthew Everett, and Pat Hanrahan. WireGL: a scalable graphics system for clusters. In ACM [ACM01], pages 129–140. ISBN 1-58113-292-1. LCCN ????

Herzog:2004:JA

[Her04] Peter Herzog. Janimation animation . . .. Computer Graphics, 38(2):8–9, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2000:CVS

[Hib00a] Bill Hibbard. Confessions of a visualization skeptic. Computer Graphics, 34(3):11–13, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2000:CGE

[Hib00b] Bill Hibbard. Contributions: from the Guest Editor. Computer Graphics, 34(1):40, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2000:VUT

[Hib00c] Bill Hibbard. Visfiles: UK team surveys Web-based visualization. Computer Graphics, 34 (1):10–12, February 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2000:VVS

[Hib00d] Bill Hibbard. Visfiles: Visual spaces. Computer Graphics, 34 (4):8–10, November 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2000:VS

[Hib00e]  Bill Hibbard.  Visualization

spaces. Computer Graphics, 34 (4):8–10, November 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2001:SIM

[Hib01a] Bill Hibbard. Super-intelligent machines. Computer Graphics, 35(1):11–13, February 2001. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2001:VFC

[Hib01b] Bill Hibbard. Vis files: computational field visualization. Computer Graphics, 35(4):5– 9, November 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2001:VFN

[Hib01c] Bill Hibbard. Vis files: NEC supercomputer visualization efforts. Computer Graphics, 35 (2):10–13, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2002:BUI

[Hib02a] Bill Hibbard. Building 3D user interface components using a visualization library. Computer Graphics, 36(1):4–7, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2002:SSC

[Hib02b] Bill Hibbard. Social synthetic characters. Computer Graphics, 36(2):5–7, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2003:G

[Hib03a] Bill Hibbard. Goodbye. Computer Graphics, 37(4):4–6, November 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2003:LOU

[Hib03b] Bill Hibbard. Large operational user of visualization. Computer Graphics, 37(3):5–9, August 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2003:SSO

[Hib03c] Bill Hibbard. Should standard oil own the roads? Computer Graphics, 37(1):5–6, February 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hibbard:2003:VBT

[Hib03d] Bill Hibbard. VisBio: a biological tool for visualization and analysis. Computer Graphics, 37(2):5–7, May 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hertzmann:2001:IA

[HJO^{+}01] Aaron Hertzmann, Charles E. Jacobs, Nuria Oliver, Brian Curless, and David H. Salesin. Image analogies. In ACM [ACM01], pages 327–340. ISBN 1-58113-292-1. LCCN ????

Hibbard:2001:VTE

[HLM01] Bill Hibbard, Bill Lorensen, and Jim Miller. Visualization toolkit extreme testing. Computer Graphics, 35(3):8–11, August 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hooks:2001:AMC

[Hoo01] Ed Hooks. Adrenaline moments and conflict: two essentials for good storytelling. Computer Graphics, 35(2):40–41, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Hilaga:2001:TMF

[HSKK01] Masaki Hilaga, Yoshihisa Shinagawa, Taku Kohmura, and Tosiyasu L. Kunii. Topology matching for fully automatic similarity estimation of 3D shapes. In ACM [ACM01], pages 203–212. ISBN 1-58113-292-1. LCCN ????

Huhn:2004:SCA

[Huh04] Carl J. Huhn. Standing at the crossroads of art & technology. Computer Graphics, 38 (2):20, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Iwata:2001:PFA

[IYNK01] Hiroo Iwata, Hiroaki Yano, Fumitaka Nakaizumi, and Ryo Kawamura. Project FEELEX: adding haptic surface to graphics. In ACM [ACM01], pages 469–476. ISBN 1-58113-292-1. LCCN ????

Jagannathan:2000:EWB

[Jag00]  Sampath Jagannathan.  Es-

cher Web-based toy. Computer Graphics, 34(2):53–54, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Jensen:2001:PBN

[JDD^{+}01] Henrik Wann Jensen, Frédo Durand, Julie Dorsey, Michael M. Stark, Peter Shirley, and Simon Premože. A physicallybased night sky model. In ACM [ACM01], pages 399–408. ISBN 1-58113-292-1. LCCN ????

Jensen:2001:PMS

[JMLH01] Henrik Wann Jensen, Stephen R. Marschner, Marc Levoy, and Pat Hanrahan. A practical model for subsurface light transport. In ACM [ACM01], pages 511–518. ISBN 1-58113-292-1. LCCN ????

Jones:2003:ADI

[Jon03]  Thom Jones.  AirlinX dig-

ital information system provides welcoming 3D to travelers. Computer Graphics, 37(1): 27–28, February 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Kalantari:2004:NVA

[Kal04] Bahman Kalantari. A new visual art medium: polynomiography. Computer Graphics, 38(3): 21–23, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Keller:2000:LWE

[KB00] Patrick Keller and Christian Babski. la fabrique: a Web3D electronic museum for the binary years. Computer Graphics, 34 (2):66–68, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Kobbelt:2001:FSS

[KBSS01] Leif P. Kobbelt, Mario Botsch, Ulrich Schwanecke, and Hans-Peter Seidel. Feature sensitive surface extraction from volume data. In ACM [ACM01], pages 57–66. ISBN 1-58113-292-1. LCCN ????

Kurt:2008:RBU

[KC08] Murat Kurt and Muhammed Gökhan Cinsdikici. Representing BRDFs using SOMs and MANs. Computer Graphics, 42(3):2:1–2:??, August 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Kurt:2009:SBM

[KE09] Murat Kurt and Dave Edwards.

A survey of BRDF models for computer graphics. Computer Graphics, 43(2):4:1–4:??, May 2009. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Kitamura:2001:ISD

[KKYK01] Yoshifumi Kitamura, Takashige Konishi, Sumihiko Yamamoto, and Fumio Kishino. Interactive stereoscopic display for three or more users. In ACM [ACM01], pages 231–240. ISBN 1-58113-292-1. LCCN ????

Klima:2000:LLW

[Kli00] Ivan Klima. Lunar landing in Web3D. Computer Graphics, 34 (2):57–58, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Knowlton:2001:FCA

[KM01] Ken Knowlton and Carl Machover. On frustrations of collaborating with artists. Computer Graphics, 35(3):22–24, August 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Kindlmann:2004:SVS

[KNB^{+}04] Gordon Kindlmann, Richard A. Normann, Arun Badi, Charles Keller, Greg M. Jones, and Christopher R. Johnson. Scientific visualization in small animal imaging. Computer Graphics, 38(2):4–7, May 2004. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Kolacinski:2008:EIY

[Kol08] Brad Kolacinski. Education: If you animate it, they will come. Computer Graphics, 42(3):9:1– 9:??, August 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lang:2000:CCG

[Lan00a] Scott Lang. Chapters continue to grow. Computer Graphics, 34 (3):16–21, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lang:2000:COS

[Lan00b] Teresa Lang. Comics from the other side. Computer Graphics, 34(4):32, November 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lang:2001:COS

[Lan01] Teresa Lang. Comics from the other side. Computer Graphics, 35(3):42, August 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lang:2002:COS

[Lan02a] Teresa Lang. Comics from the other side. Computer Graphics, 36(1):40, February 2002. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lang:2002:COSa

[Lan02b] Teresa Lang. Comics from the other side. Computer Graphics, 36(2):48, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lang:2002:COSb

[Lan02c] Teresa Lang. Comics from the other side. Computer Graphics, 36(3):48, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Laporte:2000:AGL

[Lap00] Cati Laporte. Avatar, get a life. Computer Graphics, 34 (2):61–62, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Larnder:2006:APC

[Lar06]  Chris Isaac Larnder.  Aug-

mented perception via cartoon rendering: reflections on a realtime video-to-cartoon system. Computer Graphics, 40(3):8:1– 8:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lasseter:2001:TAC

[Las01] John Lasseter. Tricks to animating characters with a computer. Computer Graphics, 35 (2):45–47, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Legakis:2001:FBC

[LDG01] Justin Legakis, Julie Dorsey, and Steven Gortler. Feature-based cellular texturing for architectural models. In ACM [ACM01], pages 309–316. ISBN 1-58113-292-1. LCCN ????

Leker:2002:EEC

[Lek02] Terri Leker. End-to-end content creation with desktop production studios. Computer Graphics, 36 (2):23–26, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lemay:2001:ALS

[Lem01] Brian Lemay. Ancient layout scrolls found in archeological dig: can these old principles be applied to the advanced technology of tomorrow? Computer Graphics, 35(2):36, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Leo:2002:YGA

[Leo02a] Etta Di Leo. Are you getting all that you can out of ACM SIG-GRAPH? Computer Graphics, 36(2):8–9, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Leo:2002:SAS

[Leo02b] Etta Di Leo. Starting an ACM SIGGRAPH chapter. Computer Graphics, 36(4):10–11, November 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Levy:2001:CTM

[Lév01] Bruno Lévy. Constrained texture mapping for polygonal meshes. In ACM [ACM01], pages 417– 424. ISBN 1-58113-292-1. LCCN ????

Lew:2009:CDC

[Lew09] Jack Lew. Connecting the dots by crossing borders. Computer Graphics, 43(2):6:1–6:??, May 2009. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lingafelter:2007:SRD

[Lin07] Daniel Lingafelter. SIGGRAPH 2007: report from Daniel Lingafelter. Computer Graphics, 41 (4):1–2, November 2007. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Lindholm:2001:UPV

[LKM01] Erik Lindholm, Mark J. Kligard, and Henry Moreton. A userprogrammable vertex engine. In ACM [ACM01], pages 149–158. ISBN 1-58113-292-1. LCCN ????

Lum:2002:IKE

[LM02] Eric B. Lum and Kwan-Liu Ma.

Interactivity is the key to expressive visualization. Computer Graphics, 36(3):5–9, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Liu:2001:EEM

[LSZ01] Zicheng Liu, Ying Shan, and Zhengyou Zhang. Expressive expression mapping with ratio images. In ACM [ACM01], pages 271–276. ISBN 1-58113-292-1. LCCN ????

Liu:2001:SBT

[LYS01] Xinguo Liu, Yizhou Yu, and Heung-Yeung Shum. Synthesizing bidirectional texture functions for real-world surfaces. In ACM [ACM01], pages 97–106. ISBN 1-58113-292-1. LCCN ????

Ma:2004:VQE

[Ma04a] Kwan-Liu Ma. Visualization: a quickly emerging field. Computer Graphics, 38(1):4–7, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ma:2004:VS

[Ma04b] Kwan-Liu Ma. Visualization for security. Computer Graphics, 38 (4):4–6, November 2004. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ma:2005:TV

[Ma05] Kwan-Liu Ma. Teaching visualization. Computer Graphics, 39 (1):4–5, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ma:2007:CCS

[Ma07a] Kwan-Liu Ma. Creating a collaborative space to share data, visualization, and knowledge. Computer Graphics, 41(4):1– 4, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ma:2007:VNS

[Ma07b] Kwan-Liu Ma. VisFiles: the next surge of visualization research. Computer Graphics, 41 (3):3:1–3:??, August 2007. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

McCool:2001:HFB

[MAA01] Michael D. McCool, Jason Ang, and Anis Ahmad. Homomorphic factorization of BRDFs for highperformance rendering. In ACM [ACM01], pages 171–178. ISBN 1-58113-292-1. LCCN ????

Machover:2000:CGPa

[Mac00a] Carl Machover. Computer graphics pioneers: looking back… to what’s ahead. Computer Graphics, 34(1):30–32, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Machover:2000:CGPb

[Mac00b] Carl Machover. Computer graphics pioneers: Opportunities and remembrances. Computer Graphics, 34(4):19–20, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Machover:2000:GHS

[Mac00c] Carl Machover. Green and Hopgood share UK computer graphics history. Computer Graphics, 34(2):28–30, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Machover:2001:CGP

[Mac01a] Carl Machover. Computer graphics pioneers: the Giloi’s school of computer graphics. Computer Graphics, 35(4):12– 16, November 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-

Machover:2001:HV

[Mac01b] Carl Machover. The history of versaCAD. Computer Graphics, 35(1):22–23, February 2001. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Masucci:2004:QEE

[Mas04] Michael Masucci. Quantum entanglement and the EZTV online museum. Computer Graphics, 38 (3):14–17, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Matsuba:2000:VDL

[Mat00] Stephen N. Matsuba. VRML dream on-line performance. Computer Graphics, 34(2):52– 53, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

McAfee:2003:CGR

[McA03] Francis X. McAfee. Computer graphics research in the humanities at the center for electronic communication. Computer Graphics, 37(4):13–16, November 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

McCann:2000:ROV

[McC00] Mike McCann. Remotely operated vehicle (ROV) dive visualization. Computer Graphics, 34(2):51, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Megahed:2000:HMP

[Meg00] Abe Megahed. Hypercosm: making 3D programming accessible. Computer Graphics, 34 (2):47–48, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Malzbender:2001:PTM

[MGW01] Tom Malzbender, Dan Gelb, and Hans Wolters. Polynomial texture maps. In ACM [ACM01], pages 519–528. ISBN 1-58113-292-1. LCCN ????

Millar:2006:HDR

[Mil06a]  Hamish Millar.  High dy-

namic range (’HDR’) technology. Computer Graphics, 40(3):6:1– 6:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Millar:2006:VCR

[Mil06b] Hamish Millar. Videogames:

content and responsibility. Computer Graphics, 40(3):7:1–7:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

McCaslin:2000:VDN

[MMS00] Paula T. McCaslin, Philip A.

McDonald, and Edward J. Szoke. 3D visualization development at NOAA forecast systems laboratory. Computer Graphics, 34(1): 41–44, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Mnich:2000:MH

[Mni00] Georg Mnich. Mystery house.

Computer Graphics, 34(2):50– 51, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

McQueen:2000:TDC

[MP00] David M. McQueen and Charles S. Peskin. A three-dimensional computer model of the human heart for studying cardiac fluid dynamics. Computer Graphics, 34(1):56–60, February 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Milenkovic:2001:OBA

[MS01] Victor J. Milenkovic and Harald Schmidl. Optimization-based animation. In ACM [ACM01], pages 37–46. ISBN 1-58113-292-1. LCCN ????

Myszkowski:2001:PGG

[MTAS01] Karol Myszkowski, Takehiro Tawara, Hiroyuki Akamine, and Hans-Peter Seidel. Perceptionguided global illumination solution for animation rendering. In ACM [ACM01], pages 221–230. ISBN 1-58113-292-1. LCCN ????

Nelson:2001:SPP

[NER01] David Nelson, Bob Ellis, and Laurie Reinhart. SIGGRAPH public policy committee activity detailed. Computer Graphics, 35 (1):24–28, February 2001. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Oh:2001:IBM

[OCDD01] Byong Mok Oh, Max Chen, Julie Dorsey, and Frédo Durand. Image-based modeling and photo editing. In ACM [ACM01], pages 433–442. ISBN 1-58113-292-1. LCCN ????

OBrien:2001:SSP

[OCE01] James F. O’Brien, Perry R.

Cook, and Georg Essl. Synthesizing sounds from physically based motion. In ACM [ACM01], pages 529–536. ISBN 1-58113-292-1. LCCN ????

Ostromoukhov:2001:SEE

[Ost01] Victor Ostromoukhov. A simple and efficient error-diffusion algorithm. In ACM [ACM01], pages 567–572. ISBN 1-58113-292-1. LCCN ????

Owen:2006:MP

[Owe06] Scott Owen. Message from the president. Computer Graphics, 40(1):5:1–5:??, February 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Pan:2000:CGA

[PaHL00] Zhigeng Pan, Pheng ann Heng, and Rynson W. H. Lau. Computer graphics around the world: computer graphics in Hong Kong. Computer Graphics, 34 (1):15–19, February 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Park:2000:VSH

[Par00a] Tae-Ha Park. Virtual show hostess — Sunny. Computer Graphics, 34(2):68, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Parnanen:2000:IIB

[Par00b] Pekka Parnanen. Interactive 3D interface for e-business solutions. Computer Graphics, 34 (2):54–55, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Petersen:2007:ADA

[Pet07] Peter Petersen. Art in the digital age from a personal perspective. Computer Graphics, 41(4): 1–3, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Pausch:2000:AMP

[PF00] Randy Pausch and Clifton Forlines. Alice: model, paint & animate — easy-to-use interactive graphics for the Web. Computer Graphics, 34(2):42–43, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Perry:2001:KSS

[PF01] Ronald N. Perry and Sarah F. Frisken. Kizamu: a system for sculpting digital characters. In ACM [ACM01], pages 47–56. ISBN 1-58113-292-1. LCCN ????

Pauly:2001:SPP

[PG01] Mark Pauly and Markus Gross. Spectral processing of pointsampled geometry. In ACM [ACM01], pages 379–386. ISBN 1-58113-292-1. LCCN ????

Praun:2001:RTH

[PHWF01] Emil Praun, Hugues Hoppe, Matthew Webb, and Adam Finkelstein. Real-time hatching. In ACM [ACM01], page 581. ISBN 1-58113-292-1. LCCN ????

Pinney:2000:SW

[Pin00] Samanth Pinney. Streaming 3D over the Web. Computer Graphics, 34(2):59, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Pitzel:2003:AI

[Pit03] Steve Pitzel. An animator at Intel. Computer Graphics, 37(3): 18–19, August 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Parish:2001:PMC

[PM01] Yoav I. H. Parish and Pascal Müller. Procedural modeling of cities. In ACM [ACM01], pages 301–308. ISBN 1-58113-292-1. LCCN ????

Prusinkiewicz:2001:UPI

[PMKL01] Przemyslaw Prusinkiewicz, Lars Mündermann, Radoslaw Karwowski, and Brendan Lane. The use of positional information in the modeling of plants. In ACM [ACM01], pages 289–300. ISBN 1-58113-292-1. LCCN ????

Proudfoot:2001:RTP

[PMTH01] Kekoa Proudfoot, William R.

Mark, Svetoslav Tzvetkov, and Pat Hanrahan. A real-time procedural shading system for programmable graphics hardware. In ACM [ACM01], pages 159– 170. ISBN 1-58113-292-1. LCCN ????

Praun:2001:CMP

[PSS01] Emil Praun, Wim Sweldens, and Peter Schröder. Consistent mesh parameterizations. In ACM [ACM01], pages 179–184. ISBN 1-58113-292-1. LCCN ????

Pai:2001:SPI

[PvdDJ^{+}01] Dinesh K. Pai, Kees van den Doel, Doug L. James, Jochen Lang, John E. Lloyd, Joshua L. Richmond, and Som H. Yau. Scanning physical interaction behavior of 3D objects. In ACM [ACM01], pages 87–96. ISBN 1-58113-292-1. LCCN ????

Rautek:2008:VIV

[RBGV08] Peter Rautek, Stefan Bruckner, Eduard Gröller, and Ivan Viola. VisFiles: Illustrative visualization: new technology or useless tautology? Computer Graphics, 42(3):4:1–4:??, August 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rowland:2006:VIV

[RF06] Chris Rowland and Steve Flack. Visualising the invisible: visualising historic shipwrecks. Computer Graphics, 40(3):2:1– 2:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ramamoorthi:2001:ERI

[RH01a] Ravi Ramamoorthi and Pat Hanrahan. An efficient representation for irradiance environment maps. In ACM [ACM01], pages 497–500. ISBN 1-58113-292-1. LCCN ????

Ramamoorthi:2001:SPF

[RH01b] Ravi Ramamoorthi and Pat Hanrahan. A signal-processing framework for inverse rendering. In ACM [ACM01], pages 117– 128. ISBN 1-58113-292-1. LCCN ????

Rhyne:2000:ASO

[Rhy00a] Theresa-Maria Rhyne. ACM SIGGRAPH outreach to the computer games community: a report on the Game Developers Conference 2000. Computer Graphics, 34(2):34, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rhyne:2000:CBM

[Rhy00b] Theresa-Marie Rhyne. Carto BOF meets at SIGGRAPH 2000. Computer Graphics, 34(3):42– 49, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rhyne:2000:TSI

[Rhy00c] Theresa-Marie Rhyne. Two stepping information technology with visualization: a viewpoint from the U.S. EPA Scientific Visualization Center. Computer Graphics, 34(1):45–47, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rohde:2007:MMI

[Roh07] Joe Rohde. From myth to mountain: insights into virtual placemaking. Computer Graphics, 41 (3):1:1–1:??, August 2007. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rouse:2000:GGC

[Rou00a] Richard Rouse III. Gaming and graphics: Computer games, not computer movies. Computer Graphics, 34(4):5–7, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rouse:2000:GGT

[Rou00b] Richard Rouse, III. Gaming and graphics: toiling with tools. Computer Graphics, 34 (1):5–9, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rouse:2001:GVN

[Rou01a] Richard Rouse. Games on the verge of a nervous breakdown: emotional content in computer games. Computer Graphics, 35 (1):6–10, February 2001. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rouse:2001:GGC

[Rou01b] Richard Rouse, III. Gaming and graphics: the console and PC: separated at birth? Computer Graphics, 35(2):5–9, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Raga:2009:ACF

[RS09] Elizabeth A. Raga and Stuart S. Sumida. Anatomical considerations in facial motion capture. Computer Graphics, 43(2): 2:1–2:??, May 2009. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Reynolds:2001:IMU

[RSR^{+}01] Matthew Reynolds, Bernd Schoner, Joey Richards, Kelly Dobson, and Neil Gershenfeld. An immersive, multi-user, musical stage environment. In ACM [ACM01], pages 553–560. ISBN 1-58113-292-1. LCCN ????

Ruiz:2000:VE

[Rui00] Kathleen Ruiz. Virtual environments/3DWeb. Computer Graphics, 34(2):59–60, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Rothfarb:2000:SFT

[RV00] Robert Rothfarb and Philippe Van Nedervelde. SimsalaGrimm: fairy-tale themed on-line 3D community. Computer Graphics, 34(2):55–56, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sauter:2003:VNM

[Sau03] Phillip M. Sauter. VR2Go^{T M}:

a new method for virtual reality development. Computer Graphics, 37(1):19–24, February 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Schneider:2008:BSP

[Sch08a] Neil Schneider. Breaking the stereoscopic 3D paradox. Computer Graphics, 42(3):5:1–5:??, August 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Schneider:2008:DRCa

[Sch08b] Neil Schneider. Dr. Robert Cailliau, co-developer of the WWW: part one. Computer Graphics, 42 (3):6:1–6:??, August 2008. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Schneider:2008:DRCb

[Sch08c] Neil Schneider. Dr. Robert Cailliau, co-developer of the WWW: part two. Computer Graphics, 42 (3):7:1–7:??, August 2008. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Scott:2003:CVL

[Sco03a] Remington Scott. Creating virtual learning environments is much closer than we think. Computer Graphics, 37(1):26, February 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Scott:2003:SLN

[Sco03b] Remington Scott. Sparking life: notes on the performance capture sessions for the Lord of the Rings: the Two Towers. Computer Graphics, 37(4):17– 21, November 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Stoll:2001:LHP

[SEP^{+}01] Gordon Stoll, Matthew Eldridge, Dan Patterson, Art Webb, Steven Berman, Richard Levy, Chris Caywood, Milton Taveira, Stephen Hunt, and Pat Hanrahan. Lightning-2: a highperformance display subsystem for PC clusters. In ACM [ACM01], pages 141–148. ISBN 1-58113-292-1. LCCN ????

Sawhney:2001:HSC

[SGH^{+}01] Harpreet S. Sawhney, Yanlin Guo, Keith Hanna, Rakesh Kumar, Sean Adkins, and Samuel Zhou. Hybrid stereo camera: an IBR approach for synthesis of very high resolution stereoscopic image sequences. In ACM [ACM01], pages 451–460. ISBN 1-58113-292-1. LCCN ????

Shamson:2000:M

[Sha00a]  Alex Shamson.  MagicBlos-

som. Computer Graphics, 34 (2):62–63, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Shaul:2000:TM

[Sha00b] Frederick Shaul. twirl — 3D microbrowser. Computer Graphics, 34(2):65–66, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sim:2000:TAS

[Sim00a] Carrie Sim. “they all suck!”

(whereas we’ve been running too fast to do anything but exhale). Computer Graphics, 34 (2):64–65, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sims:2000:VCC

[Sim00b] Ed Sims. Virtual communicator characters. Computer Graphics, 34(2):44, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sloss:2006:BFF

[Slo06] Eric Sloss. Ben Fry finding methods of visualizing data accepts Nierenberg chair of design at Carnegie Mellon. Computer Graphics, 40(2):3:1–3:??, May 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sun:2001:AGG

[SM01] Harold C. Sun and Dimitris N.

Metaxas. Automating gait generation. In ACM [ACM01], pages 261–270. ISBN 1-58113-292-1. LCCN ????

Shi:2001:CGA

[SP01] Jiaoying Shi and Zhigeng Pan. Computer graphics around the world: computer graphics in China: an overview. Computer Graphics, 35(2):22–27, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sparling:2007:GSI

[Spa07] George A. J. Sparling. Graphics and the symbolization of ideas. Computer Graphics, 41 (2):1:1–1:??, May 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Spencer:2005:EB

[Spe05a] Stephen N. Spencer. The end … and the beginning! Computer Graphics, 39(2):2–3, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Spencer:2005:LTC

[Spe05b] Stephen N. Spencer. Learning through computer-generated visualization. Computer Graphics, 39(1):2, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Shaffer:2000:RTI

[SR00] Eric Shaffer and Daniel A. Reed. Real-time immersive performance visualization and steering. Computer Graphics, 34 (2):11–14, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sander:2001:TMP

[SSGH01] Pedro V. Sander, John Snyder, Steven J. Gortler, and Hugues Hoppe. Texture mapping progressive meshes. In ACM [ACM01], pages 409–416. ISBN 1-58113-292-1. LCCN ????

Strehlke:2000:R

[Str00] Kai Strehlke. Roomz. Computer Graphics, 34(2):49–50, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sullivan:2000:ACS

[Sul00] Karen Sullivan. About the cover:

Stirring it up. Computer Graphics, 34(4):3–4, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sullivan:2001:GE

[Sul01] Karen Sullivan. From the Guest Editor. Computer Graphics, 35 (2):28–29, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Sullivan:2002:ASD

[Sul02] Karen Sullivan. Adaptively sampled distance fields. Computer Graphics, 36(1):3, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Swinstead:2008:MMH

[Swi08] Philip Swinstead. A match made in Hollywood heaven: how video games made animation and filmmaking (a lot) easier. Computer Graphics, 42(2):3:1–3:??, May 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Talbot:2004:CFE

[Tal04] Martin Talbot. Challenges fuel excitement for work. Computer Graphics, 38(4):11–13, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Tatum:2000:AW

[Tat00] Mandee Tatum. Active Worlds. Computer Graphics, 34(2):56– 57, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Taylor:2000:PSV

[Tay00] Russell M. Taylor, II. Practical scientific visualization examples. Computer Graphics, 34(1): 74–79, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Theobalt:2004:VBP

[TCMS04] Christian Theobalt, Joel Carranza, Marcus A. Magnor, and Hans-Peter Seidel. 3D video: being part of the movie. Computer Graphics, 38(3):18–20, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Taddei:2000:AVH

[TDM00] Ugo Taddei, Olaf David, and Christian Michl. Application of VisAD in hydrological modeling and simulation. Computer Graphics, 34(1):48–50, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Teo:2007:LT

[Teo07] Trina Teo. The little things. Computer Graphics, 41(4):1– 2, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Tsingos:2001:MAV

[TFNC01] Nicolas Tsingos, Thomas Funkhouser, Addy Ngan, and Ingrid Carlbom. Modeling acoustics in virtual environments using the uniform theory of diffraction. In ACM [ACM01], pages 545–552. ISBN 1-58113-292-1. LCCN ????

Thomas:2004:GPM

[Tho04a] Mark Thomas. A graphics pipeline for making 3D as cool as text. Computer Graphics, 38 (1):22–25, February 2004. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Thompson:2004:OTI

[Tho04b] Pamela Kleibrink Thompson.

Oppor-toon-ities in animation and computer graphics. Computer Graphics, 38(3):25–27, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Thorisson:2004:AIC

[TPLD04] Kristinn R. Thórisson, Christopher Pennock, Thor List, and John DiPirro. Artificial intelligence in computer graphics: a constructionist approach. Computer Graphics, 38(1):26– 30, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Tuffy:2007:ASS

[Tuf07]  Valerie Tuffy.  ACM SIG-

GRAPH Singapore Chapter trip to Bangkok. Computer Graphics, 41(1):1–27, February 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

URL  http://portalparts.
acm.org/1230000/1229086/fm/
frontmatter.pdf.

Tupper:2001:RTD

[Tup01]  Jeff Tupper.  Reliable two-

dimensional graphing methods for mathematical formulae with two free variables. In ACM [ACM01], pages 77– 86. ISBN 1-58113-292-1. LCCN ???? URL http://

www.dgp.toronto.edu/people/
mooncake/papers/SIGGRAPH2001_
Tupper.pdf.

Turk:2001:TSS

[Tur01] Greg Turk. Texture synthesis on surfaces. In ACM [ACM01], pages 347–354. ISBN 1-58113-292-1. LCCN ????

Umble:2006:MIR

[Umb06] Eden Ashley Umble. Making it real: the future of stereoscopic 3D film technology. Computer Graphics, 40(1):3:1–3:??, February 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Ursyn:2004:CIW

[Urs04] Anna Ursyn. Creation and instruction work together. Computer Graphics, 38(4):7–10, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Varnum:2005:WBL

[Var05]  Susan Varnum.  Web-based

learning fills a need. Computer Graphics, 39(1):6–7, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

vandenDoel:2001:FPB

[vdDKP01] Kees van den Doel, Paul G.

Kry, and Dinesh K. Pai. FoleyAutomatic: physically-based sound effects for interactive simulation and animation. In ACM [ACM01], pages 537–544. ISBN 1-58113-292-1. LCCN ????

vandenHengel:2008:IBM

[vdHD08] Anton van den Hengel and Anthony Dick. Image based modelling with VideoTrace. Computer Graphics, 42(2):2:1–2:??, May 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Walsh:2000:UMM

[Wal00]  Aaron E. Walsh.  Univer-

sal media: media-rich content for bandwidth-starved devices. Computer Graphics, 34 (2):37–41, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wardin:2006:CMB

[War06] Chuck Wardin. A conference much bigger than first perceived. Computer Graphics, 40(3):4:1– 4:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wyrick:2002:PPI

[WEL02] Ben Wyrick, Bob Ellis, and Myles Losch. Public policy issues heat up, attract audience. Computer Graphics, 36(1):13– 15, February 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2000:EET

[Wes00a] Thomas G. West. Enormous eyes and tiny grasping hands. Computer Graphics, 34(3):14–15, August 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2000:IRV

[Wes00b] Thomas G. West. Images and reversals: is visualization no longer a new new thing? Computer Graphics, 34(1):13–14, February 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2000:IRW

[Wes00c] Thomas G. West. Images and reversals: When the world plague was stopped by a visual artist. Computer Graphics, 34(4):11– 12, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2000:VLP

[Wes00d] Thomas G. West. Is a visualization language possible? Computer Graphics, 34(2):15–16, May 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2000:WWP

[Wes00e] Thomas G. West. When the world plague was stopped by a digital artist. Computer Graphics, 34(4):11–12, November 2000. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2001:VTN

[Wes01] Thomas G. West. Visual thinkers and Nobel Prizes. Computer Graphics, 35(1):14–15, February 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

West:2002:TLE

[Wes02a] Thomas G. West. Thinking like Einstein on the Hokule’a: the continuing puzzle of the visual. Computer Graphics, 36(3): 10–12, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2002:MVC

[Wes02b] Hans Westman. Making visible the creative process of the artist. Computer Graphics, 36 (3):3–4, August 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2002:TMM

[Wes02c] Hans Westman. There is so much more to the SIGGRAPH conference than tired feet. Computer Graphics, 36(4):3–4, November 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2003:CGI

[Wes03a] Hans Westman. Communication and the growth of information. Computer Graphics, 37(3):3, August 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2003:CAC

[Wes03b] Hans Westman. Computer art comes of age. Computer Graphics, 37(2):3, May 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2003:CGD

[Wes03c] Hans Westman. Computer graphics: defining the literal meaning of images. Computer Graphics, 37(4):3, November 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2004:EDC

[Wes04a] Hans Westman. The exploration and discovery of computer graphics. Computer Graphics, 38 (1):3, February 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2004:FJC

[Wes04b] Hans Westman. The frustrations and joys of computer graphics. Computer Graphics, 38(3):3, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2004:PGP

[Wes04c] Hans Westman. Playing the game: preparing for a career in computer graphics. Computer Graphics, 38(4):3, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2004:WYS

[Wes04d] Hans Westman. What you see is what you get: the hidden value of abstractions. Computer Graphics, 38(2):3, May 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2005:CGW

[Wes05a] Hans Westman. Computer graphics: what’s it all about? Computer Graphics, 39(2):4, May 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2005:GTE

[Wes05b] Hans Westman. Game technology extends its reach. Computer Graphics, 39(1):3, February 2005. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2006:DCG

[Wes06a] Hans Westman. At depths with computer graphics. Computer Graphics, 40(3):1:1–1:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2006:FYWa

[Wes06b] Hans Westman. Finding your way in Computer Graphics. Computer Graphics, 40(1):1:1– 1:??, February 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2006:FYWb

[Wes06c] Hans Westman. Finding your way in computer graphics. Computer Graphics, 40(2):1:1–1:??, May 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2006:MIR

[Wes06d] Hans Westman. Making it real.

Computer Graphics, 40(1):2:1– 2:??, February 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2007:SA

[Wes07a] Hans Westman. SIGGRAPH Asia. Computer Graphics, 41(4): 1–7, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2007:I

[Wes07b] Hans Westman. This issue. Computer Graphics, 41(4):1, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2008:E

[Wes08a] Hans Westman. From the editor. Computer Graphics, 42(2): 1:1–1:??, May 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2008:EFT

[Wes08b] Hans Westman. From the editor: Facts through fiction. Computer Graphics, 42(3):1:1–1:??, August 2008. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westman:2009:BBS

[Wes09]  Hans Westman.  Building

bridges: science, the arts & technology. Computer Graphics, 43 (2):1:1–1:??, May 2009. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Walter:2001:ISP

[WFM01a] Marcelo Walter, Alain Fournier, and Daniel Menevaux. Integrating shape and pattern in mammalian models. In ACM [ACM01], pages 317–326. ISBN 1-58113-292-1. LCCN ????

Watson:2001:MPV

[WFM01b] Benjamin Watson, Alinda Friedman, and Aaron McGaffey. Measuring and predicting visual fidelity. In ACM [ACM01], pages 213–220. ISBN 1-58113-292-1. LCCN ????

Wand:2001:RBA

[WFP^{+}01] Michael Wand, Matthias Fischer, Ingmar Peter, Friedhelm Meyer auf der Heide, and Wolfgang Straßer. The randomized z-buffer algorithm: interactive rendering of highly complex scenes. In ACM [ACM01], pages 361–370. ISBN 1-58113-292-1. LCCN ????

Willstedt:2006:VES

[Wil06] Anders Willstedt. Visual effects in Sweden: the little engine that could. Computer Graphics, 40(3):3:1–3:??, November 2006. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wittlief:2007:SFA

[Wit07] Kenneth Wittlief. Stereoscopic 3D film and animation: getting it right. Computer Graphics, 41 (3):2:1–2:??, August 2007. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wei:2001:TSA

[WL01] Li-Yi Wei and Marc Levoy.

Texture synthesis over arbitrary manifold surfaces. In ACM [ACM01], pages 355–360. ISBN 1-58113-292-1. LCCN ????

Wolfe:2000:RMA

[WLS00] Rosalee Wolfe, John L. Lowther, and Ching-Kuang Shene. Rendering + modeling + animation + postprocessing = computer graphics. Computer Graphics, 34 (4):15–18, November 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Westlund:2001:AAS

[WM01] Harold B. Westlund and Gary W. Meyer. Applying appearance standards to light reflection models. In ACM [ACM01], pages 501–51. ISBN 1-58113-292-1. LCCN ????

Wakita:2003:IVW

[WM03] Akira Wakita and Fumio Matsumoto. Information visualization with Web3D: spatial visualization of human activity area and its condition. Computer Graphics, 37(3):29–33, August 2003. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wolfe:2000:ERM

[Wol00] Rosalee Wolfe. Education: Rendering + modeling + animation + postprocessing = computer graphics. Computer Graphics, 34 (4):15–18, November 2000. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Woo:2004:FAS

[Woo04] Andrew Woo. Focus on ACM SIGGRAPH Canadian chapters. Computer Graphics, 38(4):14– 17, November 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Xu:2001:PRK

[XCL^{+}01] Ying-Qing Xu, Yanyun Chen, Stephen Lin, Hua Zhong, Enhua Wu, Baining Guo, and Heung-Yeung Shum. Photorealistic rendering of knitwear using the lumislice. In ACM [ACM01], pages 391–398. ISBN 1-58113-292-1. LCCN ????

Yoo:2004:TSV

[YM04] Terry S. Yoo and Kwan-Liu Ma. Taking stock of visualization in scientific computing. Computer Graphics, 38(3):4–6, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Noh:2001:EC

[yNN01] Jun yong Noh and Ulrich Neumann. Expression cloning. In ACM [ACM01], pages 277–288. ISBN 1-58113-292-1. LCCN ????

Zdravkovic:2002:UGP

[ZCHS02] Vladan Zdravkovic, Eva Carling, Anders Hast, and Stefan Seipel. University of Gävle prepares students for entry into industry. Computer Graphics, 36 (2):27–32, May 2002. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Zen:2004:PA

[Zen04] Jen Zen. Painting in air. Computer Graphics, 38(3):7–9, August 2004. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Zhao:2007:SLZ

[Zha07]  Liming Zhao.  SIGGRAPH

2007: Liming Zhao, report from Univ. Penn student chapter. Computer Graphics, 41(4): 1–2, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-

Zirkind:2007:ADC

[Zir07] Givon Zirkind. AFIS data compression: an example of how domain specific compression algorithms can produce very high compression ratios. Computer Graphics, 41(4):1–36, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Zwicker:2001:SS

[ZPvBG01] Matthias Zwicker, Hanspeter Pfister, Jeroen van Baar, and Markus Gross. Surface splatting. In ACM [ACM01], pages 371– 378. ISBN 1-58113-292-1. LCCN ????